# Matched Filtering-Based Channel Estimation for AFDM Systems in Doubly Selective Channels

### v1

@@ Abstract
Affine frequency division multiplexing (AFDM) has recently emerged as an excellent backward-compatible 6G waveform. In this paper, an enhanced AFDM is proposed whereby the delay-Doppler (DD) coupling phase is considered. Specifically, we study matched filtering (MF) assisted channel estimation (CE) for AFDM systems in complex doubly selective channels. By deriving the complete input-output relationship, the inter-chirp-carrier interference, signal-to-interference-plus-noise ratio (SINR), and the effective SINR loss of AFDM, are investigated in discrete affine Fourier transform (DAFT) domain. Further, we look into the path ambiguity problem and show that it may lead to severe performance deterioration in fractional-delay fractional-Doppler channels. To address such a problem, we introduce an MF assisted CE scheme building upon a novel pilot arrangement across two consecutive AFDM transmissions. This allows us to sequentially estimate the parameters of each path by exploiting the separability and approximate orthogonality of different paths in the DAFT domain, thus leading to significantly reduced complexity. Furthermore, based on generalized Fibonacci search (GFS), an MF-GFS scheme is proposed to avoid significantly redundant computation, which can be extended to typical wide-band systems. Extensive simulation results indicate that the proposed schemes offer superior advantages in terms of their improved communication performance and lower complexity.
Index Terms:
Affine frequency division multiplexing (AFDM), channel estimation, doubly selective channels, matched filtering.
@@ I Introduction
Several waveforms that can adapt to high-mobility scenarios have been studied. Among many others, a representative waveform is orthogonal time-frequency space (OTFS) whose information symbols are transmitted in the delay-Doppler (DD) domain through two-dimensional (2D) orthogonal basis functions [4, 5, 6, 7]. Since each information symbol in the DD domain spans the entire time-frequency grid, OTFS is able to achieve a significant improvement in error rate performance compared to OFDM [6, 7, 8, 9, 10, 11, 12, 13]. However, OTFS requires radical change for the transceiver design and hence may not permit a seamless integration of the legacy OFDM based wireless systems.
Recently, affine frequency division multiplexing (AFDM) has emerged as an excellent backward-compatible 6G waveform for efficient and reliable high-mobility communications [13, 14, 15]. With minimum modification of OFDM, AFDM modulates the data symbols using multiple orthogonal chirp-carriers. The modulation is carried out through discrete affine Fourier transform (DAFT), enabling efficient mapping between the DAFT domain and the time domain. By appropriately tuning the chirp rate according to the Doppler profile of the channel, AFDM enables proper spreading in the time-frequency domain, thus allowing it to achieve the full diversity over the doubly selective channels. Besides, the sparsity and compactness of the channel in the DAFT domain can be exploited for reducing the pilot overhead. Special cases of AFDM include DAFT-OFDM [16] and orthogonal chirp division multiplexing (OCDM) [17], yet they may not be able to achieve the full diversity [15].
A plethora of recent works have further explored AFDM. Inspired by the index modulation (IM) for OFDM [18], IM-AFDM systems were studied in [19, 20, 21, 22] for improving the spectrum efficiency. The integration of generalized spatial modulation (GSM) and AFDM, called GSM-AFDM, was studied in [23] to design low-power and high-performance multiple-input and multiple-output (MIMO) systems. For 6G integrated sensing and communications (ISAC), several studies demonstrated the advantages of AFDM empowered ISAC [24, 25, 26] systems. To support massive machine-type high-mobility communications, AFDM was also exploited as the building waveform for sparse code multiple access (SCMA) systems in [27].
@@
Against the above background, this work investigates matched filtering (MF) assisted CE schemes in doubly selective channels for AFDM systems. We differentiate integer-delay-fractional-Doppler (IDFD) and fractional-delay-fractional-Doppler (FDFD) channels and develop advanced MF-CE schemes through the co-design of transmit pilot signals and receiver signal processing. The main contributions are summarized as follows:
  • Firstly, we adopt the input-output (I/O) relationship in the time domain, in alignment with that employed in existing OTFS-related research. We derive the corresponding I/O relationship in the DAFT domain and then present a comprehensive analysis on the inter-chirp-carrier-interference (ICCI), signal-to-interference-plus-noise ratio (SINR) and effective SINR loss.
  • Secondly, we present the first analysis of the path ambiguity problem [^1] As we will show later, (37) becomes invalid for FDFD channels, thus resulting in path ambiguity. in FDFD channels, where the inability to accurately estimate the delay and Doppler values of a path leads to significant degradation in CE performance. Motivated by this, a new pilot pattern and an MF scheme are proposed. The key idea is to use the different equivalent shifts of a path in the two AFDM symbols with different chirp rates to uniquely determine the delay and Doppler values [^2] Specifically, the same path exhibits different equivalent shifts in two AFDM symbols with different chirp rates. The path ambiguity problem can be resolved by estimating the two equivalent shifts with additional pilot overhead.. In addition, by leveraging the separability and orthogonality of different paths, our proposed MF scheme is able to eliminate the matrix inversion operation, decouple joint estimation, and narrow the search range, thus leading to significantly reduced CE complexity.
  • Thirdly, based on the generalized Fibonacci search (GFS) algorithm [35, 36, 37], we propose an MF-GFS CE scheme for FDFD channels. It is found that our proposed scheme can reduce the amount of search when estimating fractional parameters, yielding a significant reduction of redundant computations whilst outperforming the proposed MF scheme.
@@
The rest of this paper is organized as follows. Section II reviews the basic concepts of AFDM. Section III analyzes the I/O relationship, ICCI, SINR, and effective SINR loss of the AFDM system. Section IV introduces the MF CE scheme and MF-GFS CE scheme. Section V gives the simulation results. Finally, Section VI concludes this paper.
@@ II-A AFDM modulation
Assume that the number of chirp subcarriers is $N$, the subcarrier spacing is $\varDelta f$ (Hz), the bandwidth is $B=N\varDelta f$, and the duration of an AFDM symbol is $T$. The DAFT domain axis is sampled at multiple times of $\varDelta f$ to obtain a discrete grid, i.e., $\boldsymbol{\varXi}=\left\{n\varDelta f,n=0,\cdots,N-1\right\}$. In this paper, the AFDM system is illustrated in Fig. 1. For the transmitter (TX) side, consider a set of $N$ information symbols $\left\{x\left(n\right),n=0,\cdots,N-1\right\}$ from a modulation alphabet $\mathbb{A}$ of size $M_{\mathrm{mod}}$ (e.g. QAM symbols), which are arranged on the DAFT domain grid $\boldsymbol{\varXi}$. The AFDM modulator first employs inverse DAFT (IDAFT) to map $x\left(n\right)$ to $s\left(\bar{n}\right)$, as shown below:
$s\left(\bar{n}\right)=\sum_{n=0}^{N-1}{x\left(n\right)}F_{c_{1},c_{2}}\left(\bar{n},n\right),$ (1)
where $F_{c_{1},c_{2}}\left(\bar{n},n\right)=\frac{1}{\sqrt{N}}e^{j2\pi\left(c_{1}\bar{n}^{2}+c_{2}n^{2}+\frac{\bar{n}n}{N}\right)}$, $\bar{n}\in\left\{0,\cdots,N-1\right\}$ is the time index, $n\in\left\{0,\cdots,N-1\right\}$ is the chirp subcarrier index in the DAFT domain, $c_{1}$ and $c_{2}$ are DAFT parameters. Note that AFDM achieves full diversity in doubly selective channels if $c_{1}$ satisfies [15]
@@
where integer $\xi$ satisfies $0<\xi\leqslant N$. (1) can be rewritten as
@@
where $\boldsymbol{x}=\left[x\left(0\right),\cdots,x\left(N-1\right)\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{s}=\left[s\left(0\right),\cdots,s\left(N-1\right)\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{A}=\boldsymbol{\varLambda}_{c_{2}}\boldsymbol{F\varLambda}_{c_{1}}$ is the normalized DAFT matrix, $\boldsymbol{F}$ is the normalized DFT matrix, $\boldsymbol{\varLambda}_{c}=\mathrm{diag}\mathrm{(}e^{-j2\pi cn^{2}},n=0,1,...,N-1)$, $\left(\cdot\right)^{\mathrm{T}}$ is the transpose operation, $\left(\cdot\right)^{\mathrm{H}}$ is the conjugate transpose operation, $\mathrm{diag}\left(\cdot\right)$ is to convert a vector into a diagonal matrix or extract the diagonal elements of a matrix. Before transmission, a chirp-periodic prefix (CPP) [15] of length $L_{\mathrm{CPP}}$ is also employed.
@@ II-B Doubly Selective Channels
where $h_{i}$, $\tau_{i}\in\left[0,\tau_{\max}\right]$ and $\nu_{i}\in\left[-\nu_{\max},\nu_{\max}\right]$ are the channel gain, delay and Doppler parameters of the $i$-th path, respectively, $\delta\left(\cdot\right)$ is the Dirac delta function, $P$ is the number of paths. $\nu_{\max}$ and $\tau_{\max}$ are the maximum Doppler and delay, respectively. The normalized delay $l_{\tau_{i}}$ and Doppler $k_{\nu_{i}}$ satisfy
@@
[…]
@@
$r(t)=\iint{h(\tau,\nu)s(t-\tau)e^{j2\pi\nu t}\mathrm{d}\tau\mathrm{d}\nu}+\tilde{w}\left(t\right),$ (6)
@@
Substitute (4) into (7), and consider $r\left(t\right)$ which is sampled at multiples of $T_{s}=\small{\frac{1}{N\varDelta f}}=\small{\frac{T}{N}}$. After discarding CPP, the vector $\boldsymbol{r}=\left[r\left(0\right),\cdots,r\left(N-1\right)\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$ is
$r(\bar{n})=\sum_{i=1}^{P}{h_{i}s(\bar{n}-l_{\tau_{i}})e^{\frac{j2\pi k_{\nu_{i}}}{N}\left(\bar{n}-l_{\tau_{i}}\right)}}+\tilde{w}\left(\bar{n}\right).$ (8)
where the noise $\tilde{w}\left(\bar{n}\right)$ follows $\mathcal{C}\mathcal{N}\left(\tilde{w}\left(\bar{n}\right)\left|0,\lambda^{-1}\right.\right)$, $\lambda$ is the noise accuracy. (8) can be expressed in matrix form as
$\boldsymbol{r}=\boldsymbol{H}_{\mathrm{T}}\boldsymbol{s}+\tilde{\boldsymbol{w}},$ (9)
where $\boldsymbol{H}_{\mathrm{T}}=\sum_{i=1}^{P}{h_{i}}\boldsymbol{\varGamma}_{\mathrm{CPP}_{i}}\boldsymbol{\varPi}^{l_{\tau_{i}}}\boldsymbol{\varDelta}^{k_{\nu_{i}}}\in\mathbb{C}^{N\times N}$ is the channel matrix in the time domain, $\boldsymbol{\varDelta}=\mathrm{diag}\left(z^{0},z^{1},\cdots z^{N-1}\right)\in\mathbb{C}^{N\times N}$, $z=e^{\frac{j2\pi}{N}}$, $\boldsymbol{\varPi}\in\mathbb{C}^{N\times N}$ is the forward cyclic-shift matrix, $\boldsymbol{\varGamma}_{\mathrm{CPP}_{i}}\in\mathbb{C}^{N\times N}$ is used to model CPP [15]. The delay and Doppler values are modeled by $\boldsymbol{\varPi}$ and $\boldsymbol{\varDelta}$, respectively.

### v3

@@ Abstract
Affine frequency division multiplexing (AFDM) has recently emerged as an excellent backward-compatible 6G waveform. In this paper, we study matched filtering (MF) assisted channel estimation (CE) for AFDM systems in complex doubly selective channels. By deriving the complete input-output relationship of the continuous-time signal, the inter-chirp-carrier interference, signal-to-interference-plus-noise ratio (SINR), and the effective SINR loss of AFDM, are investigated in discrete affine Fourier transform (DAFT) domain. Further, we propose two low-complexity methods for constructing the channel matrix by taking advantage of its inherent discrete Fourier transform structure and the staircase structure of the piecewise functions in the channel matrix, respectively. It is shown that complexity reduction by at least two orders of magnitude can be achieved for a large number of chirp subcarriers. For the CE problem in doubly selective channels, we introduce an MF assisted CE scheme. This allows us to sequentially estimate the parameters of each path by exploiting the separability and approximate orthogonality of different paths in the DAFT domain, thus leading to significantly reduced complexity. Furthermore, based on generalized Fibonacci search (GFS), an MF-GFS scheme is proposed to avoid significantly redundant computation, which can be extended to typical wide-band systems. Extensive simulation results indicate that the proposed schemes offer superior advantages in terms of their improved communication performance and lower complexity.
@@ I Introduction
Several waveforms that can adapt to high-mobility scenarios have been studied. Among many others, a representative waveform is orthogonal time-frequency space (OTFS) whose information symbols are transmitted in the delay-Doppler (DD) domain through two-dimensional (2D) orthogonal basis functions [4, 5, 6]. Since each information symbol in the DD domain spans the entire time-frequency grid, OTFS is able to achieve a significant improvement in error rate performance compared to OFDM [5, 6, 7, 8, 9, 10, 11]. In addition, Zak-OTFS [12, 13] has emerged as a strong competitor to conventional OTFS in recent years, offering an alternative Zak-transform-based formulation that preserves the DD domain advantages while enabling flexible waveform implementations. However, OTFS requires radical change for the transceiver design and hence may not permit a seamless integration into the legacy OFDM based wireless systems.
Recently, affine frequency division multiplexing (AFDM) has emerged as an excellent backward-compatible 6G waveform for efficient and reliable high-mobility communications [14, 15, 16, 17, 18, 19, 20, 21]. With minimum modification of OFDM, AFDM modulates the data symbols using multiple orthogonal chirp-carriers. The modulation is carried out through discrete affine Fourier transform (DAFT), enabling efficient mapping between the DAFT domain and the time domain. By appropriately tuning the chirp rate according to the Doppler profile of the channel, AFDM enables proper spreading in the time-frequency domain, thus allowing it to achieve the full diversity over the doubly selective channels. Besides, the sparsity and compactness of the channel in the DAFT domain can be exploited for reducing the pilot overhead. Special cases of AFDM include DAFT-OFDM [22] and orthogonal chirp division multiplexing (OCDM) [23], yet they may not be able to achieve the full diversity [16].
A plethora of recent works have further explored AFDM. By flexibly selecting the chirp parameter, AFDM can achieve significant performance improvements in several aspects. Inspired by the index modulation (IM) for OFDM [24], IM-AFDM systems were studied in [25, 26, 27, 28, 29] for improving the spectrum efficiency. From the PAPR reduction aspect, the chirp-permuted AFDM was employed in [29], while the grouped pre-chirp selection algorithm was proposed in [30].
The integration of generalized spatial modulation (GSM) and AFDM, called GSM-AFDM, was studied in [31] to design low-power and high-performance multiple-input and multiple-output (MIMO) systems. For 6G integrated sensing and communications (ISAC), several studies demonstrated the advantages of AFDM empowered ISAC [32, 33, 34, 35] systems. To support massive machine-type high-mobility communications, AFDM was also exploited as the building waveform for sparse code multiple access (SCMA) systems in [36].
@@
Based on the correlation between AFDM basis functions and the received signal, this work investigates matched filtering (MF) assisted CE in doubly selective channels for AFDM systems. We differentiate integer-delay-fractional-Doppler (IDFD) and fractional-delay-fractional-Doppler (FDFD) channels and develop advanced MF-CE schemes. The main contributions are summarized as follows:
  • Firstly, we adopt the input-output (I/O) relationship in the time domain, in alignment with that employed in existing OTFS-related research. We derive the corresponding I/O relationship for continuous-time signals and then present a comprehensive analysis on the inter-chirp-carrier-interference (ICCI), signal-to-interference-plus-noise ratio (SINR) and effective SINR loss.
  • Secondly, for FDFD channels, we propose two low-complexity methods for fast channel matrix construction. The first method exploits the inherent DFT structure of the channel matrix, enabling efficient computation via fast fourier transform (FFT). The second method leverages the staircase structure of the piecewise functions in the channel matrix to directly calculate partial sums. For large subcarrier numbers, the proposed methods can reduce the complexity by at least two orders of magnitude.
  • Thirdly, for FDFD channels, we propose an MF CE scheme with joint DD estimation (MF-JE). By leveraging the separability and orthogonality of different paths, our proposed MF scheme is able to eliminate the matrix inversion operation, decouple multipath estimation, and narrow the search range, thus leading to significantly reduced complexity. In addition, by decoupling the DD estimation to further narrow the search region, a more efficient MF (MF-DE) CE scheme is developed at the cost of a slight performance loss.
  • Fourthly, based on the generalized Fibonacci search (GFS) algorithm [^1] GFS is an unconstrained nonlinear optimization method for unimodal functions. The ratio of two consecutive generalized Fibonacci numbers (GFNs) approximates the golden ratio. In GFS, two consecutive GFNs are used to non-uniformly divide the search interval, thereby efficiently narrowing the search range, reducing complexity, and enabling more accurate estimation. [42, 43, 44], we propose an MF-GFS-DE CE scheme for FDFD channels. It is found that our proposed scheme can reduce the amount of search when estimating fractional parameters, yielding a significant reduction of redundant computations whilst outperforming the proposed MF scheme.
@@
The rest of this paper is organized as follows. Section II reviews the basic concepts of AFDM. Section III analyzes channel in the DAFT domain. Section IV introduces the MF CE scheme and MF-GFS CE scheme. Section V gives the simulation results. Finally, Section VI concludes this paper.
Notations: The $m$-th element of vector $\boldsymbol{x}$ is denoted by $x\left[m\right]$, the element in the $m$-th row and $n$-th column of matrix $\boldsymbol{X}$ is $X\left[m,n\right]$. $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ denote the conjugate, transpose, and Hermitian operations, respectively. $\mathrm{diag}\left(\cdot\right)$ is to convert a vector into a diagonal matrix or extract the diagonal elements of a matrix. $\lceil\cdot\rceil$, $\lfloor\cdot\rfloor$, and $\lfloor\cdot\rceil$ denote the ceiling, floor, and nearest-integer operators, respectively. $E\left\{\cdot\right\}$ is the expected operation. $\left\{\boldsymbol{x}\times\boldsymbol{y}\right\}$ is the Cartesian product. $\delta\left(\cdot\right)$ is the Dirac delta function. $\boldsymbol{I}_{N}$ is the $N\times N$ identity matrix. $\left|\cdot\right|$ denotes the modulus operation. $\left[\cdot\right]_{N}$ is the mod-$N$ operation.
@@ II-A AFDM modulation
Assume that the number of chirp subcarriers is $N$, the subcarrier spacing is $\varDelta f$ (Hz), the bandwidth is $B=N\varDelta f$, and the duration of an AFDM symbol is $T$. The DAFT domain axis is sampled at multiple times of $\varDelta f$ to obtain a discrete grid, i.e., $\boldsymbol{\varXi}=\left\{n\varDelta f,n=0,\cdots,N-1\right\}$. In this paper, the AFDM system is illustrated in Fig. 1. For the transmitter (TX) side, consider a set of $N$ information symbols $\left\{x\left[n\right],n=0,\cdots,N-1\right\}$ from a modulation alphabet $\mathbb{A}$ of size $M_{\mathrm{mod}}$ (e.g. QAM symbols), which are arranged on the DAFT domain grid $\boldsymbol{\varXi}$. The AFDM modulator first employs inverse DAFT (IDAFT) to map $x\left[n\right]$ to $s\left[\bar{n}\right]$, as shown below:
$s\left[\bar{n}\right]=\sum_{n=0}^{N-1}{x\left[n\right]}\phi_{n}\left[\bar{n}\right],$ (1)
where $\phi_{n}\left[\bar{n}\right]=\frac{1}{\sqrt{N}}e^{j2\pi\left(c_{1}\bar{n}^{2}+c_{2}n^{2}+\frac{\bar{n}n}{N}\right)}$, $\bar{n}\in\left\{0,\cdots,N-1\right\}$ is the time index, $n\in\left\{0,\cdots,N-1\right\}$ is the chirp subcarrier index in the DAFT domain, $c_{1}$ and $c_{2}$ are DAFT parameters. Note that with a proper $c_{1}$, the paths are separable in the DAFT domain, enabling AFDM to achieve full diversity in doubly selective channels, i.e., [16]
@@
where $\xi$ is a small non-negative integer introduced to suppress fractional Doppler. This paper sets $c_{1}$ to the value in (2) by default. Other values of $c_{1}$ may be possible as long as AFDM achieves full diversity. (1) can be rewritten as
@@
where $\boldsymbol{x}=\left[x\left[0\right],\cdots,x\left[N-1\right]\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{s}=\left[s\left[0\right],\cdots,s\left[N-1\right]\right]^{\mathrm{T}}\in\mathbb{C}^{N\times 1}$, $\boldsymbol{A}=\boldsymbol{\varLambda}_{c_{2}}\boldsymbol{F\varLambda}_{c_{1}}$ is the normalized DAFT matrix, $\boldsymbol{F}$ is the normalized DFT matrix, $\boldsymbol{\varLambda}_{c}=\mathrm{diag}\mathrm{(}e^{-j2\pi cn^{2}},n=0,1,...,N-1)$. Before transmission, a chirp-periodic prefix (CPP) [16] of length $L_{\mathrm{CPP}}$ is also employed. To reflect the spectral wrapping inherent in the AFDM signal, the continuous transmit signal in $0\leqslant t<T$ should be expressed as [21]
$s\left(t\right)=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}{x\left[n\right]g_{\mathrm{tx}}\left(t\right)\phi_{n}\left(t\right),}$ (4)
where $g_{\mathrm{tx}}\left(t\right)$ is the transmit pulse-shaping filter, $\phi_{n}\left(t\right)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{T_{s}^{2}}t^{2}+\frac{n}{T}t+\varphi_{n}\left(t\right)\right)}$, $T_{s}=\frac{T}{N}$ is the sampling period, $\varphi_{n}\left(t\right)=\sum_{q=0}^{2Nc_{1}}{\alpha_{n,q}\left(t\right)}$, $q\in\left\{0,1,\cdots,2Nc_{1}\right\}$, $\alpha_{n,q}\left(t\right)$ is given by
$\alpha_{n,q}\left(t\right)=\begin{cases}\lfloor\frac{q\left(q-\frac{n}{N}\right)}{2c_{1}}\rfloor-\frac{q}{T_{s}}t,&t_{n,q}\leq t<t_{n,q+1}\\ 0,&\mathrm{otherwise}\\ \end{cases},$ (5)
$t_{n,q}=\begin{cases}0,&q=0\\ \frac{qN-n}{2Nc_{1}}T_{s},&1\leq q\leq 2Nc_{1}\\ \end{cases}.$ (6)
@@ II-B Doubly Selective Channels
where $h_{i}$, $\tau_{i}\in\left[0,\tau_{\max}\right]$ and $\nu_{i}\in\left[-\nu_{\max},\nu_{\max}\right]$ are the channel gain, delay and Doppler parameters of the $i$-th path, respectively, $P$ is the number of paths. $\nu_{\max}$ and $\tau_{\max}$ are the maximum Doppler and delay, respectively. The normalized delay $l_{\tau_{i}}$ and Doppler $k_{\nu_{i}}$ satisfy
@@
where $k_{\nu_{i}}=\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}\in\left[-k_{\max},k_{\max}\right]$, $l_{\tau_{i}}=\bar{l}_{\tau_{i}}+\iota_{\tau_{i}}\in\left[0,l_{\max}\right]$, $\bar{k}_{\nu_{i}}$ and $\bar{l}_{\tau_{i}}$ are both integer parts, $\kappa_{\nu_{i}}\in\left[-0.5,0.5\right]$ and $\iota_{\tau_{i}}\in\left[-0.5,0.5\right]$ are both fractional parts, $k_{\max}=\lceil\frac{N\nu_{\max}}{B}\rceil$, $l_{\max}=\lceil\tau_{\max}B\rceil$. Due to the finite $B$ and $T$, there is generally fractional delay and Doppler. $\nu_{\max}$ and $\tau_{\max}$ can usually be considered known a priori [9]. However, the number of scatterers in the environment is usually unpredictable, resulting in the uncertainty of $P$.
@@
$r(t)=\iint{h(\tau,\nu)s(t-\tau)e^{j2\pi\nu t}\mathrm{d}\tau\mathrm{d}\nu}+\tilde{w}\left(t\right).$ (9)
@@
After discarding CPP, $r\left(t\right)$ is given by
$r(t)$ $=\sum_{i=1}^{P}{h_{i}s\left(t-\tau_{i}\right)e^{j2\pi\nu_{i}\left(t-\tau_{i}\right)}}+\tilde{w}\left(\bar{n}\right),$ (11)
where $\tilde{w}\left(\bar{n}\right)$ is a zero-mean Gaussian noise with variance $\sigma^{2}$.